OpenStudy (anonymous):

I have the answer can someone just see if I am correct? Rewrite the following function in the form f(x)=kx^p. f(x)=(2x)(5square root x) k= p=

OpenStudy (anonymous):

\[f(x)=(2x)(5\sqrt{x})\]

OpenStudy (jtvatsim):

what do you have?

OpenStudy (anonymous):

k=10 and p= 2

OpenStudy (jtvatsim):

k is good, how did you get p?

OpenStudy (anonymous):

well that is were I'm not sure if I did it right what do you do with the square root

OpenStudy (jtvatsim):

ok, the trick is \[\sqrt{x}=x^{\frac{1}{2}}\]

OpenStudy (anonymous):

oh so if there are no numbers always =.5

OpenStudy (anonymous):

so then k=10 and p=.5 correct

OpenStudy (jtvatsim):

close, you have one x with nothing so that is x^1 and another x with x^(1/2) so you get x^(1 + 1/2) = x^(3/2)

OpenStudy (anonymous):

k= 10 and then p=1.5 right

OpenStudy (jtvatsim):

to clarify what you said earlier if there are no numbers AND there is a square root than it is .5 :)

OpenStudy (jtvatsim):

correct!

OpenStudy (anonymous):

ok thank you

OpenStudy (jtvatsim):

the general formula is this: \[\sqrt[a]{x^b}=x^{\frac{b}{a}}\]

OpenStudy (anonymous):

ok

OpenStudy (jtvatsim):

no problem!